OpenStudy (anonymous):

a triangle has sides : 8 1/3, 5 4/5, 9 3/10 inches and a height of 4 3/4 inches find the (a) the perimeter and (b) the area

OpenStudy (anonymous):

if the sides of the triangle are a,b,c then perimeter =a+b+c and area =sqrt(s*(s-a)*(s-b)*(s-c)) where s=(a+b+c)/2

OpenStudy (anonymous):

thank you (: